OpenStudy (anonymous):

If a^1/n=r,then which of the following are true statements? Check all that apply. A.r^n=a B. n√a = r C. ar = n D. n1/r = a

OpenStudy (anonymous):

\[B.\sqrt[n]{a}=r\]

OpenStudy (anonymous):

@shifuyanli @SolomonZelman Help D:<

OpenStudy (anonymous):

roots

OpenStudy (anonymous):

D.

OpenStudy (solomonzelman):

can you draw the options or write them in latex in equation editor so that I can see what the options (precisely) are?

OpenStudy (solomonzelman):

If you want to get me feedback....

OpenStudy (anonymous):

D.

OpenStudy (solomonzelman):

@my feedback

OpenStudy (anonymous):

\[A.r^{n}=a\] \[B.\sqrt[n]{a}=r\] \[C.a^r=n\] \[D. n^{1/r}=a\]

OpenStudy (anonymous):

@SolomonZelman sorry first time using open study >//<

OpenStudy (solomonzelman):

B is the initial statement, right? that is what you are started off, so it is certainly not derived, but it is however a true statement. A is what you get when you raise both sides to the power of n. So it is true. C is not (necessarily) true. it is false. D is false too.

OpenStudy (solomonzelman):

as far as the latex goes you can use `\(\Large\color{black}{ }\)` For example when you say `\(\Large\color{black}{ \sqrt[n]{a}=r }\)`, you will get \(\Large\color{black}{ \sqrt[n]{a}=r }\) you can make it blue also, `\(\Large\color{blue}{ \sqrt[n]{a}=r }\)` this gives, \(\Large\color{blue}{ \sqrt[n]{a}=r }\).

OpenStudy (solomonzelman):

like you can play around switch sizes and colors.

OpenStudy (solomonzelman):

you can do fonts boxes and anything... but this is just for now...

OpenStudy (solomonzelman):

specific question message me or anyone when he/she is online. If a user doesn't know it's fine, but if he does, you will get your help....

OpenStudy (solomonzelman):

bye:)

OpenStudy (anonymous):

Thanks!!!!!!!!!!!!!!! @SolomonZelman

OpenStudy (solomonzelman):

Anytime:) \(\huge\color{ goldenrod }{\huge {\bbox[5pt, lightyellow ,border:2px solid lightyellow ]{ ☀ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\\ \large\color{black}{ \bf \large\color{ black }{~~~~~~~\large {\bbox[5pt, red ,border:2px solid black ]{ {~^{\large⊞}~\huge{~▓ }~~\large^{⊞} ~ } }~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~}} } \\ \color{lightyellow}{d~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~} }}}\)

OpenStudy (solomonzelman):

oh, question marks, it is supposed to be a house when symbols are seen correctly, but that ain't matters.

OpenStudy (solomonzelman):

byw:)

OpenStudy (anonymous):

xD totally your biggest fan now !

OpenStudy (anonymous):

what about me?